# SpikeSTAG: Spatial-Temporal Forecasting via GNN-SNN Collaboration

### v1

@@ Abstract
Introduction
Spiking Neural Networks (SNNs), recognized as the third generation of neural networks, excel in processing temporal data by virtue of their precise temporal coding and intrinsically low energy consumption(?). Consequently, they have been widely adopted for tasks including image classification (?????), visual processing (??), sequential image recognition (??), and time-series classification (??). Owing to their ability to encode and transmit information through discrete spikes, SNNs exhibit exceptional spatial-temporal modeling capabilities, achieving performance comparable to or exceeding that of traditional Artificial Neural Networks (ANNs) on short-sequence tasks. For long-sequence modeling, the potential of SNNs is being progressively unlocked through novel architectures(??).
Although prior studies have showcased the remarkable temporal modeling capabilities of SNNs, a truly general-purpose architecture often needs both spatial and temporal reasoning (???) Current SNNs, however, still underperform conventional networks in capturing rich spatial structures. Recent efforts therefore seek to equip SNNs with stronger spatial inductive biases(??), thereby broadening their applicability across diverse domains.
Multivariate time-series forecasting leverages intricate inter-variable dependencies to anticipate future dynamics, thereby empowering decision-makers to optimize resource allocation, mitigate risk, and devise more effective strategies. Its applications span finance, energy, transportation, and beyond, consistently enhancing the precision and efficiency of operational decisions. Reliable forecasts demand simultaneous modeling of temporal evolution and spatial interactions; substantial research has been devoted to this dual objective. One line of work enhances predictive accuracy by coupling convolutional neural networks (CNNs) with recurrent architectures, most notably through CNN–LSTM hybrids (???). Another line of research integrates recurrent networks or attention mechanisms into deep learning frameworks to further improve temporal modeling (??). Alternatively, graph neural networks provide a distinct paradigm by explicitly modeling spatial dependencies among variables (??).
@@ Introduction
By converting continuous-valued time series into meaningful spike trains, recent work has pioneered a training paradigm that, for the first time, enables SNNs to tackle multivariate forecasting, unleashing their inherent temporal modeling prowess (?). Yet this seminal study—along with most follow-ups—largely overlooks spatial dependencies. Subsequent efforts have concentrated solely on refining the temporal inductive biases of SNNs (?). Consequently, all existing models privilege temporal optimization and neglect spatial representation, leaving a critical void in the joint modeling of spatial and temporal patterns for multivariate time-series prediction.
@@
Related Work
GNNs have demonstrated strong potential in time-series tasks—forecasting, classification, anomaly detection, and imputation—whenever data exhibit spatial or relational structure (?). Numerous studies cast time series as graphs of interacting segments or motifs, leveraging GNNs to capture relational patterns and enable personalized search or recommendation. These methods, however, rely on predefined topologies, constraining their capacity to learn adaptive node connections and generalize across domains (??). In forecasting tasks, models such as STGCN, MTGNN, and AGCRN explicitly construct graphs—e.g., sensor networks in traffic prediction—to encode spatial dependencies among variables (???). Yet their spatial-temporal modules remain loosely coupled, limiting the interaction between spatial and temporal representations and consequently impeding predictive performance.
@@
On the other hand, SNNs, as biologically inspired models, exhibit strong potential in temporal sequence modeling. Recent efforts—including SpikeTCN, SpikeRNN, and Spike-Transformer—have extended SNNs to time-series forecasting. These models first transform continuous signals into spike trains via dedicated encoders such as Delta or convolutional coders, and then leverage SNN neurons to capture temporal dynamics (?),. However, SNNs still face challenges in effectively integrating spatial and temporal capabilities, particularly in complex multivariate time series tasks, where both dimensions are critical for performance.
In recent years, SNNs have been increasingly investigated for graph-structured data (???). Existing methods typically apply Poisson encoding to transform node features into spike trains and then perform graph message passing via SNN neurons, demonstrating viability on static-graph tasks such as node classification and link prediction. To address dynamic graphs, subsequent work has extended SNNs to temporal domains, exploiting the sparsity of spike encoding and spike-driven attention to enhance efficiency and scalability (??). Nevertheless, these approaches primarily emphasize structural exploration of the graph itself, neglecting the SNN’s inherent ability to model continuous-time dynamics.
While existing works innovatively introduce SNNs to graph domains for tasks like node classification and link prediction, they fail to leverage SNNs’ temporal dynamics for enhanced time modeling. Instead, SNNs merely replace traditional neurons in GNNs as computational units, resulting in decoupled spatial-temporal modeling. Furthermore, these approaches are restricted to structural inference tasks (e.g., node/graph classification) and overlook spatial-temporally coupled applications such as traffic flow or power load forecasting—neglecting SNNs’ inherent advantages in temporal processing.
@@
Multi-variate Time-series Forecasting
@@
Spiking Neural Networks
In an SNN, one of the fundamental processing units is the Leaky Integrate-and-Fire (LIF) neuron (?), which models membrane potential as the charge stored on a capacitor. The neuronal dynamics are governed by three sequential operations: integration, firing, and reset.
[…]
@@
To address the non-differentiability of the Heaviside step function $\Theta(\cdot)$ during back-propagation, we adopt a surrogate-gradient method that replaces the exact derivative with the derivative of an arctangent function $\sigma^{\prime}(x)=\frac{\alpha}{2}\left(1+\left(\frac{\pi}{2}\alpha x\right)^{2}\right)$, which exhibits a sigmoid-like shape and provides a biased but effective gradient estimator for end-to-end training (?).
Method
@@
Model Architecture
SpikeSTAG is a spatial-temporal architecture that unites spiking neural computation with graph learning. Raw node features are first fused with auxiliary temporal features (minute-of-hour, hour-of-day, day-of-week) to form an enriched embedding; from this embedding, learnable node parameters induce an adaptive adjacency matrix. An Observation (OBS) Block refines each node’s representation across time, after which Multi-scale Spike Aggregation (MSSA) gathers fine-grained neighborhood information and uses spikes to aggregate. Finally, a Dual-path Spike Fusion (DSF) layer translates the spike output into a sequence of hidden states via a lightweight LSTM, which is then discretized via spike encoding. The resulting spike trains are processed by the Spike Self-Attention (SSA) mechanism, yielding membrane potentials that are used to generate the final prediction.
@@
Embedded Series and Adaptive Matrix
@@
In virtually all forecasting scenarios, temporal cues are indispensable for grasping the global context, and the most salient indicators revolve around “where” and “when”. Therefore, during temporal embedding, we enrich each node’s feature vector with dataset-specific side information—variables, hour-of-day, and day-of-week.
@@
Learning an adjacency matrix directly from node features is non-trivial. In SpikeSTAG, we adopt the widely-used yet elegant source-target attention mechanism to build an adaptive graph (?). The similarity between source node $i$ and target node $j$ is measured by an inner product and normalized into a link probability via the sigmoid function:
@@
where $\mathbf{E}\in\mathbf{R}^{N\times d}$ denote the learnable node embeddings; each row $\mathbf{e}_{i}$ represents the latent feature vector of node $i$ after fusing the raw node attributes with auxiliary temporal covariates. $\sigma(\cdot)$ is the element-wise sigmoid that converts unbounded similarities into link probabilities (0,1$\in$(0,1)∈ ( 0 , 1 ), ensuring a fully-adaptive, weighted adjacency. To further guarantee that a node’s own history remains the dominant contributor to its future state, we explicitly incorporate self-loops into the resulting matrix.
Observation Block
Figure 3 illustrates the Observation (OBS) Block, which refines node features that have already been embedded along the temporal dimension to yield time-augmented representations. However, at this stage, the node information remains insufficiently integrated; we therefore introduce a graph-attention mechanism to aggregate neighborhood information and learn the interrelations among temporally embedded nodes across different timestamps. In contrast to standard attention, we restrict the computation to relationships within each node’s local neighborhood; the formula details are shown in Appendix A, where $x_{i}^{\prime}$ represents the enhanced model.
@@
Multi-Scale Spiking Aggregation
@@
The Multi-Scale Spiking Aggregation (MSSA) module is proposed to refine the conventional GraphSAGE framework(?), which consists of two key components: importance-based sampling and spiking aggregation, as illustrated in Figure 2. Traditional layer-wise sampling directly feeds the raw adjacency matrix into the model, inevitably injecting noise and distracting the model from genuine relational signals. To suppress the interference of low-weight edges, we first prune the outgoing edges of node $i$ via a learnable threshold, yielding a candidate set $C_{i}$ composed of salient neighbors only.
@@
[…]
@@
where $\mathbf{x}_{j}\!\in\!\{0,1\}^{T}$ is the spike train of node $j$. This vector is then passed through an LIF layer. The resulting sparse spike trains are forwarded to downstream SNN or recurrent cells; the identical sample–index–sum–LIF routine is repeated for the two-hop neighborhood $\mathcal{S}_{i}^{(2)}$. Because the sampling budget $|\mathcal{S}_{i}^{(k)}|\!\ll\!N$, both memory footprint and runtime scale linearly with graph size, rendering the entire MSSA module entirely free of dense GEMM and conferring substantial empirical speed-ups.
@@
Dual-Path Spike Fusion
@@
Datasets
Table 1 summarizes the benchmark datasets employed. Long-sequence datasets consist of Electricity (hourly power consumption in kWh) and Solar (photovoltaic generation logs). Short-sequence datasets comprise METR-LA (mean traffic speed on Los Angeles County highways) and PEMS-BAY (mean traffic speed in the San Francisco Bay Area). For datasets whose sampling granularity is one minute, the auxiliary-node features comprise three signals—minute-of-hour, hour-of-day, and day-of-week. Conversely, for datasets sampled at hourly intervals, only hour-of-day and day-of-week are provided. As evaluation metrics, we adopt the root relative squared error (RSE) and the coefficient of determination ($R^{2}$). Owing to their invariance to the absolute numerical scale of the data, these two measures are preferred over mean squared error (MSE) or mean absolute error (MAE) in time-series forecasting studies.
Main Results
@@
SpikeSTAG’s explicit spatial modeling yields robust generalization across domains. On spatially correlated data, the model surpasses iSpikformer notably. For the PEMS-BAY traffic network, SpikeSTAG records an $RSE$ of 0.661 at the 24-step horizon, 0.6 percent lower than iSpikformer’s 0.665. When the horizon extends to 48 steps, SpikeSTAG raises its $R^{2}$ to 0.403, exceeding iSpikformer’s 0.348 by 15.8 percent. These outcomes confirm SpikeSTAG as the state of the art among spiking temporal forecasting networks.
@@
Hyper-parameter Sensitivity
As an SNN-based architecture, SpikeSTAG treats the time-step $T_{s}$ as a hyper-parameter that dictates the granularity at which temporal dynamics $\Delta t$ are modeled. Experiments were conducted on METR-LA with a prediction horizon $L=3$ while varying $T_{s}\in\{4,8,12,16\}$. Figures 4(a)–4(b) reveal that a modest increase in the number of time steps yields a slight rise in the $R^{2}$ index because finer temporal resolution becomes achievable, yet the overall impact on network performance remains negligible; SpikeSTAG thereby demonstrates remarkable stability.
When $T_{s}=16$, $R^{2}$ exhibits a marginal decline. This observation aligns with the “self-accumulating dynamics”(?), in which surrogate-gradient-induced error accumulation can precipitate vanishing or exploding gradients and consequently degrade model efficacy. The complementary $RSE$ metric displays an inverse trend. Consequently, in most experimental settings, a smaller $T_{s}$ can be adopted to further reduce computational energy without sacrificing accuracy.
Figure 4: The impact of the time step $T_{s}$, a key hyperparameter, in SNN: (a) shows the $R^{2}$ values for different time steps, and (b) shows the $RSE$ values for different time steps.
Ablation Study

### v2

@@ Abstract
1 Introduction
Spiking Neural Networks (SNNs), recognized as the third generation of neural networks, excel in processing temporal data by virtue of their precise temporal coding and intrinsically low energy consumptionmaass1997networks . Consequently, they have been widely adopted for tasks including image classification hu2021spiking ; fang2021deep ; ding2021optimal ; zhou2022spikformer ; yao2023spike , visual processing meftah2010segmentation ; escobar2009action , sequential image recognition jeffares2021spike ; chen2023unleashing , and time-series classification dominguez2018deep ; fang2020multivariate . Owing to their ability to encode and transmit information through discrete spikes, SNNs exhibit exceptional spatial-temporal modeling capabilities, achieving performance comparable to or exceeding that of traditional Artificial Neural Networks (ANNs) on short-sequence tasks. For long-sequence modeling, the potential of SNNs is being progressively unlocked through novel architecturesshen2025spikingssms ; zhong2024spikessm .
Although prior studies have showcased the remarkable temporal modeling capabilities of SNNs, a truly general-purpose architecture often needs both spatial and temporal reasoning ain2016structural ; yan2018spatial ; musa2023deep Current SNNs, however, still underperform conventional networks in capturing rich spatial structures. Recent efforts therefore seek to equip SNNs with stronger spatial inductive biaseszhang2025staa ; yu2022stsc , thereby broadening their applicability across diverse domains.
Multivariate time-series forecasting leverages intricate inter-variable dependencies to anticipate future dynamics, thereby empowering decision-makers to optimize resource allocation, mitigate risk, and devise more effective strategies. Its applications span finance, energy, transportation, and beyond, consistently enhancing the precision and efficiency of operational decisions. Reliable forecasts demand simultaneous modeling of temporal evolution and spatial interactions; substantial research has been devoted to this dual objective. One line of work enhances predictive accuracy by coupling convolutional neural networks (CNNs) with recurrent architectures, most notably through CNNâLSTM hybrids jin2019prediction ; livieris2020cnn ; widiputra2021multivariate . Another line of research integrates recurrent networks or attention mechanisms into deep learning frameworks to further improve temporal modeling zhang1998time ; liu2023itransformer . Alternatively, graph neural networks provide a distinct paradigm by explicitly modeling spatial dependencies among variables yi2023fouriergnn ; cao2020spectral .
@@ 1 Introduction
By converting continuous-valued time series into meaningful spike trains, recent work has pioneered a training paradigm that, for the first time, enables SNNs to tackle multivariate forecasting, unleashing their inherent temporal modeling prowess lv2024efficient . Yet this seminal studyâalong with most follow-upsâlargely overlooks spatial dependencies. Subsequent efforts have concentrated solely on refining the temporal inductive biases of SNNs shibo2025tslif . Consequently, all existing models privilege temporal optimization and neglect spatial representation, leaving a critical void in the joint modeling of spatial and temporal patterns for multivariate time-series prediction.
@@
2 Related Work
GNNs have demonstrated strong potential in time-series tasksâforecasting, classification, anomaly detection, and imputationâwhenever data exhibit spatial or relational structure jin2024survey . Numerous studies cast time series as graphs of interacting segments or motifs, leveraging GNNs to capture relational patterns and enable personalized search or recommendation. These methods, however, rely on predefined topologies, constraining their capacity to learn adaptive node connections and generalize across domains yang2019aligraph ; pal2020pinnersage . In forecasting tasks, models such as STGCN, MTGNN, and AGCRN explicitly construct graphsâe.g., sensor networks in traffic predictionâto encode spatial dependencies among variables han2020stgcn ; wu2020connecting ; bai2020adaptive . Yet their spatial-temporal modules remain loosely coupled, limiting the interaction between spatial and temporal representations and consequently impeding predictive performance.
@@ 2 Related Work
On the other hand, SNNs, as biologically inspired models, exhibit strong potential in temporal sequence modeling. Recent effortsâincluding SpikeTCN, SpikeRNN, and Spike-Transformerâhave extended SNNs to time-series forecasting. These models first transform continuous signals into spike trains via dedicated encoders such as Delta or convolutional coders, and then leverage SNN neurons to capture temporal dynamics lv2024efficient ,. However, SNNs still face challenges in effectively integrating spatial and temporal capabilities, particularly in complex multivariate time series tasks, where both dimensions are critical for performance.
In recent years, SNNs have been increasingly investigated for graph-structured data dold2021spike ; chian2021learning ; xu2021exploiting . Existing methods typically apply Poisson encoding to transform node features into spike trains and then perform graph message passing via SNN neurons, demonstrating viability on static-graph tasks such as node classification and link prediction. To address dynamic graphs, subsequent work has extended SNNs to temporal domains, exploiting the sparsity of spike encoding and spike-driven attention to enhance efficiency and scalability li2023scaling ; sun2024spikegraphormer . Nevertheless, these approaches primarily emphasize structural exploration of the graph itself, neglecting the SNNâs inherent ability to model continuous-time dynamics.
While existing works innovatively introduce SNNs to graph domains for tasks like node classification and link prediction, they fail to leverage SNNs’ temporal dynamics for enhanced time modeling. Instead, SNNs merely replace traditional neurons in GNNs as computational units, resulting in decoupled spatial-temporal modeling. Furthermore, these approaches are restricted to structural inference tasks (e.g., node/graph classification) and overlook spatial-temporally coupled applications such as traffic flow or power load forecastingâneglecting SNNs’ inherent advantages in temporal processing.
@@ 3 Preliminary
3.1 Multi-variate Time-series Forecasting
@@
3.2 Spiking Neural Networks
In an SNN, one of the fundamental processing units is the Leaky Integrate-and-Fire (LIF) neuron gerstner2002spiking , which models membrane potential as the charge stored on a capacitor. The neuronal dynamics are governed by three sequential operations: integration, firing, and reset.
$\left\{\begin{aligned} U[t]&=I[t]+H[t-\Delta t],\\ S[t]&=\Theta\!\bigl{(}U[t]-U_{\mathrm{th}}\bigr{)},\\ H[t]&=\beta\,U[t]\bigl{(}1-S[t]\bigr{)}+U_{\mathrm{reset}}S[t].\end{aligned}\right$. (1)
@@ 3.2 Spiking Neural Networks
To address the non-differentiability of the Heaviside step function $\Theta(\cdot)$ during back-propagation, we adopt a surrogate-gradient method that replaces the exact derivative with the derivative of an arctangent function $\sigma^{\prime}(x)=\frac{\alpha}{2}\left(1+\left(\frac{\pi}{2}\alpha x\right)^{2}\right)$, which exhibits a sigmoid-like shape and provides a biased but effective gradient estimator for end-to-end training fang2021deep .
4 Method
@@
4.1 Model Architecture
SpikeSTAG is a spatial-temporal architecture that unites spiking neural computation with graph learning. Raw node features are first fused with auxiliary temporal features (minute-of-hour, hour-of-day, day-of-week) to form an enriched embedding; from this embedding, learnable node parameters induce an adaptive adjacency matrix. An Observation (OBS) Block refines each nodeâs representation across time, after which Multi-scale Spike Aggregation (MSSA) gathers fine-grained neighborhood information and uses spikes to aggregate. Finally, a Dual-path Spike Fusion (DSF) layer translates the spike output into a sequence of hidden states via a lightweight LSTM, which is then discretized via spike encoding. The resulting spike trains are processed by the Spike Self-Attention (SSA) mechanism, yielding membrane potentials that are used to generate the final prediction.
@@ 4.1 Model Architecture
4.2 Embedded Series and Adaptive Matrix
@@
In virtually all forecasting scenarios, temporal cues are indispensable for grasping the global context, and the most salient indicators revolve around âwhereâ and âwhenâ. Therefore, during temporal embedding, we enrich each nodeâs feature vector with dataset-specific side informationâvariables, hour-of-day, and day-of-week.
@@
Learning an adjacency matrix directly from node features is non-trivial. In SpikeSTAG, we adopt the widely-used yet elegant source-target attention mechanism to build an adaptive graph wu2019graphwavenet . The similarity between source node $i$ and target node $j$ is measured by an inner product and normalized into a link probability via the sigmoid function:
@@
where $\mathbf{E}\in\mathbf{R}^{N\times d}$ denote the learnable node embeddings; each row $\mathbf{e}_{i}$ represents the latent feature vector of node $i$ after fusing the raw node attributes with auxiliary temporal covariates. $\sigma(\cdot)$ is the element-wise sigmoid that converts unbounded similarities into link probabilities $\in(0,1)$, ensuring a fully-adaptive, weighted adjacency. To further guarantee that a nodeâs own history remains the dominant contributor to its future state, we explicitly incorporate self-loops into the resulting matrix.
4.3 Observation Block
Figure 3 illustrates the Observation (OBS) Block, which refines node features that have already been embedded along the temporal dimension to yield time-augmented representations. However, at this stage, the node information remains insufficiently integrated; we therefore introduce a graph-attention mechanism to aggregate neighborhood information and learn the interrelations among temporally embedded nodes across different timestamps. In contrast to standard attention, we restrict the computation to relationships within each nodeâs local neighborhood; the formula details are shown in Appendix A, where $x_{i}^{\prime}$ represents the enhanced model.
@@ 4.3 Observation Block
4.4 Multi-Scale Spiking Aggregation
@@
The Multi-Scale Spiking Aggregation (MSSA) module is proposed to refine the conventional GraphSAGE frameworkhamilton2017inductive , which consists of two key components: importance-based sampling and spiking aggregation, as illustrated in Figure 2. Traditional layer-wise sampling directly feeds the raw adjacency matrix into the model, inevitably injecting noise and distracting the model from genuine relational signals. To suppress the interference of low-weight edges, we first prune the outgoing edges of node $i$ via a learnable threshold, yielding a candidate set $C_{i}$ composed of salient neighbors only.
@@
Subsequently, we perform a perceptual-importance truncation $\mathrm{Imp}_{i}=\sum_{j=1}^{N}A_{i,j}$, selecting the top-$k$ highest-ranked nodes to form the neighborhood set for the current sampling layer. To further enlarge the receptive field with minimal computational overhead, we devise a two-level sampling paradigmââlocalâ followed by âsemi-globalââwhich efficiently acquires a global receptive scope while maintaining sampling efficiency.
@@
where $\mathbf{x}_{j}\!\in\!\{0,1\}^{T}$ is the spike train of node $j$. This vector is then passed through an LIF layer. The resulting sparse spike trains are forwarded to downstream SNN or recurrent cells; the identical sampleâindexâsumâLIF routine is repeated for the two-hop neighborhood $\mathcal{S}_{i}^{(2)}$. Because the sampling budget $|\mathcal{S}_{i}^{(k)}|\!\ll\!N$, both memory footprint and runtime scale linearly with graph size, rendering the entire MSSA module entirely free of dense GEMM and conferring substantial empirical speed-ups.
@@
4.5 Dual-Path Spike Fusion
@@ 5 Experiments
5.1 Datasets
Table 1 summarizes the benchmark datasets employed. Long-sequence datasets consist of Electricity (hourly power consumption in kWh) and Solar (photovoltaic generation logs). Short-sequence datasets comprise METR-LA (mean traffic speed on Los Angeles County highways) and PEMS-BAY (mean traffic speed in the San Francisco Bay Area). For datasets whose sampling granularity is one minute, the auxiliary-node features comprise three signalsâminute-of-hour, hour-of-day, and day-of-week. Conversely, for datasets sampled at hourly intervals, only hour-of-day and day-of-week are provided. As evaluation metrics, we adopt the root relative squared error (RSE) and the coefficient of determination ($R^{2}$). Owing to their invariance to the absolute numerical scale of the data, these two measures are preferred over mean squared error (MSE) or mean absolute error (MAE) in time-series forecasting studies.
5.2 Main Results
@@
SpikeSTAGâs explicit spatial modeling yields robust generalization across domains. On spatially correlated data, the model surpasses iSpikformer notably. For the PEMS-BAY traffic network, SpikeSTAG records an $RSE$ of 0.661 at the 24-step horizon, 0.6 percent lower than iSpikformerâs 0.665. When the horizon extends to 48 steps, SpikeSTAG raises its $R^{2}$ to 0.403, exceeding iSpikformerâs 0.348 by 15.8 percent. These outcomes confirm SpikeSTAG as the state of the art among spiking temporal forecasting networks.
@@
5.3 Hyper-parameter Sensitivity
As an SNN-based architecture, SpikeSTAG treats the time-step $T_{s}$ as a hyper-parameter that dictates the granularity at which temporal dynamics $\Delta t$ are modeled. Experiments were conducted on METR-LA with a prediction horizon $L=3$ while varying $T_{s}\in\{4,8,12,16\}$. Figures 4(a)â4(b) reveal that a modest increase in the number of time steps yields a slight rise in the $R^{2}$ index because finer temporal resolution becomes achievable, yet the overall impact on network performance remains negligible; SpikeSTAG thereby demonstrates remarkable stability.
When $T_{s}=16$, $R^{2}$ exhibits a marginal decline. This observation aligns with the âself-accumulating dynamicsâfang2020multivariate , in which surrogate-gradient-induced error accumulation can precipitate vanishing or exploding gradients and consequently degrade model efficacy. The complementary $RSE$ metric displays an inverse trend. Consequently, in most experimental settings, a smaller $T_{s}$ can be adopted to further reduce computational energy without sacrificing accuracy.
Figure 4: The impact of the time step $T_{s}$, a key hyperparameter, in SNN: (a) shows the $R^{2}$ values for different time steps, and (b) shows the RSERSE values for different time steps.
5.4 Ablation Study